QUESTION 2 Using the data below, perform the calculations necessary to evaluate the V/Q matching for oxygen, and select either correct or incorrect for the statements below for this 50 year old male who is breathing 25 times/minute. Pbar 690 mm Hg Inspired 02%-30% PaO2 50 mm Hg PaCO2 36 mm Hg SaO2 75% Hb 12 g/dl # The PAO2 is approximately 148 mm Hg A. B. Incorrect # The patient is hypoxemic The subject is hyperventilating # The a/A-0.34 The PaO2/FIO2 ratio 1.66Please explain your answer and indicate whether each option is correct or incorrect.
Answer #1

Answer:-

The Ventilation-Perfusion (V/Q) proportion

The ventilation-perfusion proportion is - the proportion between the measure of air getting to the alveoli (the alveolar ventilation, V, in ml/min) and the measure of blood being sent to the lungs (the cardiovascular yield or Q - likewise in ml/min). Figuring the V/Q proportion is very simple -

V/Q = alveolar ventilation/cardiovascular yield

V/Q = (4 l/min)/(5 l/min) =

V/Q=0.8

1 .The PAO2 is approximately 148 mm Hg - Incorrect (approx. 106.25)

2. The patient is hypoxemic – Incorrect

3. The subject is hyperventilating - Correct (to a lesser extent)

4. The a/A = 0.34 - Incorrect- higher

5. The PaO2/FIO2 ratio = 1.66 - Incorrect - high

please do rate answer.............